# B-STaR: Monitoring and Balancing Exploration and Exploitation in Self-Taught Reasoners

### v1

@@ 1 Introduction
The experimental results from mathematical problem-solving, coding challenges and commonsense reasoning demonstrate that B-STaR significantly surpasses other self-improvement methods through balanced exploration and exploitation. For instance, B-STaR achieves a significant improvement in Pass@1 on both GSM8K and MATH, surpassing various self-improving variants while maintaining a steady upward trajectory, as depicted in Figure LABEL:fig:main_pass@1. Furthermore, we demonstrate that exploration-related metrics, such as Pass@32, are continuously improving without any notable degradation as in other baselines.
@@ 2.1 Background: Self-Improvement
(3) Improving: The selected dataset is then used to update $P_{t-1}$, producing $P_{t}$. To differentiate the generation model $M$ from the reward model, $M$ is also referred to as the policy model following RL literature. In reasoning tasks that we are going to focus on in this paper, SFT loss is commonly used in the Improving step due to its robustness and scalability (Pang et al., 2024; Dubey et al., 2024), as more sophisticated RL losses can be unstable to optimize and scale up. When SFT loss is adopted, the Rewarding step aims to reject low-quality responses and use the remaining high-quality data for training, thus this process is also referred to as rejection fine-tuning (RFT, Yuan et al. (2023)). In Appendix C, we provide detailed explanations of self-improvement concepts, including reward functions and RFT.
@@
The iterative procedure described above can be contextualized within the reinforcement learning framework (Singh et al., 2023), and the iterative design shifts the vanilla offline training towards a more dynamic online variant – when iteration intervals are short and the optimizer is inherited between iterations, the training essentially transforms into a fully online learning algorithm. Iterative RFT implementations in previous works typically adopt long iteration intervals where each iteration processes all available queries (Zelikman et al., 2022; Sun et al., 2024). Sometimes, these implementations opted to restart the training from the initial checkpoint rather than from the last saved model (Zelikman et al., 2022; Singh et al., 2023). However, we argue that always starting from the beginning is not scalable for large datasets, particularly in a streaming setting, instead, a continuous training approach more aligned with RL principles is preferable. Transitioning from offline to online training, online RFT (Shao et al., 2024) has demonstrated its superiority over traditional offline RFT methods by switching iterations more frequently, ensuring that the synthetic data remains on-policy. In this study, we explore online RFT as our primary framework for self-improvement.
@@ 2.2 The Critical Factors – Exploration and Exploitation
In this work, we mainly focus on complex problem-solving tasks such as math and coding domains, where the correctness of the responses can be easily verified on labeled datasets. [^2] Strictly speaking, the response may contain incorrect steps even though the final answer is correct, we do not further distinguish this difference following others (Zelikman et al., 2022; Singh et al., 2023) as it is not the focus of this work. This property facilitates the quantification of exploration and exploitation, for which we detail the metrics below:
@@ 3.2 Dynamic or Static – On the Configurations of Exploration and Exploitation
Exploration – temperature.
3.3 B-STaR
Building on the findings from §3.2 that the optimal configurations to maximize balance score is dynamically changing, we propose B-STaR, short for Balanced Self-Taught Reasoners, a method that maximizes the average balance score by dynamically adjusting configurations to balance exploration and exploitation. We also include related work on dynamically adjusting configurations in the Appendix F. Specifically, we adjust temperature and threshold automatically at every iteration, to maximize the average balance score. Notably, we only need to compute the balance score on a small subset of training queries to decide the balanced configurations, thus incurring negligible additional costs compared to the baselines. For example, we only use 600 MATH queries in our experiments. Our B-STaR algorithm is illustrated in Figure 4 and summarized in Algorithm 1. There are other configurations affecting balance score, such as the number of responses drawn per query (sample size) that will influence the exploration. In our preliminary trials in Figure LABEL:fig:sample, larger sample size tends to be generally helpful. Therefore, we just use the maximum sample size according to our computing budget and fix it, and mainly focus on dynamic adjustments of temperature and threshold in this work.
Methods GSM 8K MATH APPS ARC-C
P@1 P@32 P@32-4 P@1 P@32 P@32-4 P@1 P@32 P@32-4 P@1
SFT 36.6 88.5 62.2 17.0 60.8 31.2 9.3 43.5 25.5
Rest-EM (w/o RM) 40.5 89.9 69.8 22.8 60.0 33.6 14.5 43.9 28.2 70.7
Rest-EM (w/ RM) 46.3 90.7 72.2 24.2 62.8 37.4
Iterative RFT (w/o RM) 42.8 88.9 71.3 24.2 63.4 38.2 15.2 44.3 28.0 70.3
Iterative RFT (w/ RM) 46.6 90.2 74.9 24.4 62.6 39.0
Online RFT (w/o RM) 44.0 88.1 69.7 23.0 57.2 38.2 17.3 45.8 27.8 71.2
Online RFT (w/ RM) 46.8 91.4 76.5 23.2 62.6 39.2
B-STaR 53.8 93.6 81.0 27.8 67.2 42.2 19.6 49.3 30.7 73.0
Step 500 1000 1500 2000 2500 3000 3500 4000 4500
Temperature 0.5 0.8 0.9 1 1.1 1.1 0.9 1.1 1.1
Reward threshold 0 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1
Balance Score 0.470 0.538 0.589 0.621 0.646 0.660 0.673 0.678 0.679
4 Main Experiments
4.1 Setup
We evaluate B-STaR’s effectiveness in enhancing self-improvement for mathematical problem-solving, coding challenges and commonsense reasoning. Our evaluation compares several baseline methods: STaR/ResT-EM  (Zelikman et al., 2022; Singh et al., 2023), Iterative RFT, and Online RFT (Shao et al., 2024). The STaR/ResT-EM approach involves multiple iterations, where each iteration samples from the latest policy model but resets and retrains the model from scratch. In contrast, Iterative RFT builds on each previous iteration by inheriting the checkpoint and resuming training from that point. Online RFT goes a step further by not only inheriting the checkpoint but also the optimizer state and learning rate schedule, ensuring a smoother transition across iterations. Additionally, we implement these methods with two types of reward function: "without RM" (Answer) and "with RM" (Answer+PRM), as described in §2.3. We also provide detailed experimental setup in Appendix  B.
For mathematical problem-solving, we use the Mistral-7B model as our base and conduct SFT on the MATH training dataset for three epochs. The first epoch serves as the starting point for the model’s self-improvement phase. During this phase, we use Pass@1, Pass@K-S, and Reward@K-S metrics to track changes in the performance and model’s exploration and exploitation capabilities. On coding challenges, we follow  Singh et al. (2023) and adopt the APPS (Hendrycks et al., 2021) dataset for both training and testing. We use Llama-3-8B (Dubey et al., 2024) as our base model, [^3] We do not use Mistral-7B as in the math domain because we found it performs poorly on coding tasks. keeping the rest of the settings consistent with those of the math domain. We do not apply reward models to the coding task and instead use unit tests as the binary reward, which means only the sampling temperature is automatically adjusted in B-STaR.
Methods GSM 8K MATH
Online RFT 46.8 23.2
B-STaR (Temperature Adjustment Only) 53.1 25.0
B-STaR (Reward Threshold Adjustment Only) 49.1 24.6
B-STaR (Temperature + Reward Threshold) 53.8 27.8
Methods GSM8K MATH APPS ARC-C
SFT 49.4 18.8 15.6 78.8
Rest-EM (w/RM) 60.2 28.2 16.4 85.5
Iterative RFT (w/RM) 55.3 27.2 17.1 85.2
Online RFT (w/RM) 59.7 27.8 16.9 85.2
B-STaR 61.6 29.2 18.1 86.3
4.2 Results
As shown in Figure LABEL:fig:main_pass@1, all the self-improvement methods exhibit significant performance improvement after the first iteration. However, as the number of iterations increases, the growth trends of other baseline methods slows down and eventually stagnate. In contrast, B-STaR maintains a substantial growth rate and consistently outperforms other baselines. This suggests that balancing exploration and exploitation is crucial for achieving stable and efficient self-improvement, further validating the generalizability of B-STaR.
4.3 Ablation Study and Additional Results
Ablation on Dynamic Adjustments.
Generalizing to More Powerful Models.
5 Discussion
In this paper, we reveal the importance of the balance between exploration and exploitation in self-improving training, and we propose B-STaR, a novel algorithm to strike a better balance and achieve superior performance on mathematical reasoning, coding, and commonsense reasoning. While this paper adopts a simple method that manipulates exploration and exploitation through hyperparameter configurations, we expect more flexible control of these two factors can be achieved in the future to strike a better balance, such as advanced decoding approaches to directly control the exploration of the generated data, and reward model update to improve exploitation.
Appendix A Experiments for the Case Study
A.1 Experiment Setup
A.2 Results on GSM8K
Step 500 1000 1500 2000 2500 3000 3500 4000 4500
Temperature 0.65 0.75 1.05 0.95 1.05 0.85 1.05 1.15 1.05
Balance Score 0.500 0.557 0.591 0.626 0.652 0.665 0.679 0.682 0.684
@@ Appendix B Experiment Setup for Main Experiments
B.1 mathematical problem-solving

### v2

@@ 1 Introduction
The experimental results from mathematical problem-solving, coding challenges and commonsense reasoning demonstrate that B-STaR significantly surpasses other self-improvement methods through balanced exploration and exploitation. For instance, B-STaR achieves a significant improvement in Pass@1 on both GSM8K and MATH, surpassing various self-improving variants while maintaining a steady upward trajectory, as depicted in Figure LABEL:fig:main_pass@1. Furthermore, we demonstrate that exploration-related metrics, such as Pass@32, are continuously improving without any notable degradation.
@@ 2.1 Background: Self-Improvement
(3) Improving: The selected dataset is used to update $P_{t-1}$, producing $P_{t}$. To differentiate the generation model $M$ from the reward model, $M$ is also referred to as the policy model following RL literature. In reasoning tasks that we are going to focus on, SFT loss is commonly used in the Improving step due to its robustness and scalability (Pang et al., 2024; Dubey et al., 2024), as more sophisticated RL losses can be unstable to optimize and scale up. When SFT loss is adopted, the Rewarding step aims to reject low-quality responses and use remaining high-quality data for training, thus this process is also referred to as rejection fine-tuning (RFT, Yuan et al. (2023)). In Appendix C, we provide detailed explanations of self-improvement, including reward functions and RFT.
@@
The iterative procedure described above can be contextualized within the reinforcement learning framework (Singh et al., 2023), and the iterative design shifts the vanilla offline training towards a more dynamic online variant – when iteration intervals are short and the optimizer is inherited between iterations, the training essentially transforms into a fully online learning algorithm. Iterative RFT implementations in previous works typically adopt long iteration intervals where each iteration processes all available queries (Zelikman et al., 2022; Sun et al., 2024). Sometimes, these implementations opted to restart the training from the initial checkpoint rather than from the last saved model (Zelikman et al., 2022; Singh et al., 2023). However, we argue that always starting from the beginning is not scalable for large datasets, particularly in a streaming setting, instead, a continuous training approach more aligned with RL principles is preferable. Transitioning from offline to online training, online RFT (Shao et al., 2024) has demonstrated its superiority over traditional offline RFT methods by switching iterations more frequently, ensuring that the synthetic data remains on-policy. In this study, we explore online RFT as our primary framework.
@@ 2.2 The Critical Factors – Exploration and Exploitation
In this work, we mainly focus on complex problem-solving tasks such as math and coding domains, where the correctness of the responses can be easily verified on labeled datasets. [^2] Strictly speaking, the response may contain incorrect steps even though the final answer is correct, we do not further distinguish this difference following others (Zelikman et al., 2022) as it is not the focus of this work. This property facilitates the quantification of exploration and exploitation, for which we detail the metrics below:
@@ 3.2 Dynamic or Static – On the Configurations of Exploration and Exploitation
Exploration – temperature.
3.3 B-STaR
Building on the findings from §3.2 that the optimal configurations to maximize balance score is dynamically changing, we propose B-STaR, short for Balanced Self-Taught Reasoners, a method that maximizes the average balance score by dynamically adjusting configurations to balance exploration and exploitation. We also include related work on dynamically adjusting configurations in the Appendix F. Specifically, we adjust temperature and threshold automatically at every iteration, to maximize the average balance score. Notably, we only need to compute the balance score on a small subset of training queries to decide the balanced configurations, thus incurring negligible additional costs compared to the baselines. For example, we only use 600 MATH queries in our experiments. Our B-STaR algorithm is illustrated in Figure 4 and summarized in Algorithm 1. There are other configurations affecting balance score, such as the number of responses drawn per query (sample size) that will influence the exploration. In our preliminary trials in Figure LABEL:fig:sample, larger sample size tends to be generally helpful. Therefore, we just use the maximum sample size according to our computing budget and fix it, and mainly focus on dynamic adjustments of temperature and threshold.
Methods GSM 8K MATH APPS ARC-C
P@1 P@32 P@32-4 P@1 P@32 P@32-4 P@1 P@32 P@32-4 P@1
SFT 36.6 88.5 62.2 17.0 60.8 31.2 9.3 43.5 25.5
Rest-EM (w/o RM) 40.5 89.9 69.8 22.8 60.0 33.6 14.5 43.9 28.2 70.7
Rest-EM (w/ RM) 46.3 90.7 72.2 24.2 62.8 37.4
Iterative RFT (w/o RM) 42.8 88.9 71.3 24.2 63.4 38.2 15.2 44.3 28.0 70.3
Iterative RFT (w/ RM) 46.6 90.2 74.9 24.4 62.6 39.0
Online RFT (w/o RM) 44.0 88.1 69.7 23.0 57.2 38.2 17.3 45.8 27.8 71.2
Online RFT (w/ RM) 46.8 91.4 76.5 23.2 62.6 39.2
B-STaR 53.8 93.6 81.0 27.8 67.2 42.2 19.6 49.3 30.7 73.0
4 Main Experiments
4.1 Setup
We evaluate B-STaR’s effectiveness in enhancing self-improvement for mathematical problem-solving, coding challenges and commonsense reasoning. Our evaluation compares several baseline methods: STaR/ResT-EM  (Zelikman et al., 2022; Singh et al., 2023), Iterative RFT, and Online RFT (Shao et al., 2024). Additionally, we implement these methods with two types of reward function: "without RM" (Answer) and "with RM" (Answer+PRM), as described in §2.3. We also provide detailed baseline methods and experimental setup in Appendix  B.
4.2 Results
As shown in Figure LABEL:fig:main_pass@1, all the self-improvement methods exhibit significant performance improvement after the first iteration. However, as the number of iterations increases, the growth trends of other baseline methods slows down and eventually stagnate. In contrast, B-STaR maintains a substantial growth rate and consistently outperforms other baselines. This suggests that balancing exploration and exploitation is crucial for achieving stable and efficient self-improvement.
Step 500 1000 1500 2000 2500 3000 3500 4000 4500
Temperature 0.5 0.8 0.9 1 1.1 1.1 0.9 1.1 1.1
Reward threshold 0 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1 -0.1
Balance Score 0.470 0.538 0.589 0.621 0.646 0.660 0.673 0.678 0.679
4.3 Ablation Study and Additional Results
Methods GSM 8K MATH
Online RFT 46.8 23.2
B-STaR (Temperature Adjustment Only) 53.1 25.0
B-STaR (Reward Threshold Adjustment Only) 49.1 24.6
B-STaR (Temperature + Reward Threshold) 53.8 27.8
Methods GSM8K MATH APPS ARC-C
SFT 49.4 18.8 15.6 78.8
Rest-EM (w/RM) 60.2 28.2 16.4 85.5
Iterative RFT (w/RM) 55.3 27.2 17.1 85.2
Online RFT (w/RM) 59.7 27.8 16.9 85.2
B-STaR 61.6 29.2 18.1 86.3
Ablation on Dynamic Adjustments.
Generalizing to More Powerful Models.
5 Discussion
In this paper, we reveal the importance of the balance between exploration and exploitation in self-improving training, and we propose B-STaR, a novel algorithm to strike a better balance and achieve superior performance on reasoning and coding tasks. While we adopt a simple method that manipulates exploration and exploitation through hyperparameter configurations, we expect more flexible control of these two factors can be achieved in the future to strike a better balance, such as advanced decoding approaches to directly control the exploration of the generated data, and reward model update to improve exploitation.
Appendix A Experiments for the Case Study
A.1 Experiment Setup
A.2 Results on GSM8K
Step 500 1000 1500 2000 2500 3000 3500 4000 4500
Temperature 0.65 0.75 1.05 0.95 1.05 0.85 1.05 1.15 1.05
Balance Score 0.500 0.557 0.591 0.626 0.652 0.665 0.679 0.682 0.684
@@ Appendix B Experiment Setup for Main Experiments
B.1 Baseline
Our evaluation compares several baseline methods: STaR/ResT-EM  (Zelikman et al., 2022; Singh et al., 2023), Iterative RFT, and Online RFT (Shao et al., 2024). The STaR/ResT-EM approach involves multiple iterations, where each iteration samples from the latest policy model but resets and retrains the model from scratch. In contrast, Iterative RFT builds on each previous iteration by inheriting the checkpoint and resuming training from that point. Online RFT goes a step further by not only inheriting the checkpoint but also the optimizer state and learning rate schedule, ensuring a smoother transition across iterations.
B.2 mathematical problem-solving